# label*label*footnotetext: Corresponding authors

### v1

@@ 3 Rate of Convergence
Proof: From Lemma 2, we have
@@
$\leq$ $\dfrac{1}{1+x^{2}}\left[1+x^{2}+\dfrac{x^{2}}{\beta(\beta+1)}+\dfrac{x}{\beta n}\right]$
@@
Taking supremum over $x\in[0,\infty)$ in the above inequality, gives the result. $\blacksquare$
@@
Theorem 1
Let $\beta>1$, $f\in E$ and $\omega_{B+1}(f,\delta)$, $(B>0)$ be its modulus of continuity on the finite interval $[0,B+1]\subset[0,\infty)$, then
@@
where $\delta_{n}(\beta,B)=\dfrac{B}{n\beta}$ and $M_{f}(\beta,B)$ is an absolute constant depending on $f$, $\beta$ and $B$.
Proof: Let $\beta>0$ be fixed. For $x\in[0,B]$ and $t\leq B+1$, we have following well-known inequality
@@
where $\delta>0$. Now, for $x\in[0,B]$ and $t>B+1$, using the fact that $t-x>1$, we have
@@
$\leq$ $6A_{f}(1+B^{2})\left(\dfrac{B}{n\beta}\right)+\left(1+\dfrac{\left[\dfrac{B}{n\beta}\right]^{1/2}}{\delta}\right)\omega_{B+1}(f,\delta)$
@@
where $M_{f}(\beta,B)=6A_{f}(1+B^{2})$ and $\delta_{n}(\beta,B)=\dfrac{B}{n\beta}$. Hence the proof. $\blacksquare$
@@ 4 $A$-statistical Convergence
provided the series converges for each $n$. $A$ is said to be regular if $\lim_{n}(Ax)_{n}=L$ whenever $\lim_{n}x_{n}=L$. We say that, the sequence $x=(x_{n})$ is $A$-statistically convergent to $L$ and write $st_{A}-\lim_{n}x_{n}=L$ if for every $\epsilon>0$, $\lim_{n}\sum_{k:|x_{k}-L|\geq\epsilon}a_{nk}=0$.
Replacing $A$ by $C_{1}$, the Cesáro matrix of order one, the $A$-statistical convergence reduces to statistical convergence. Similarly, if we take $A=I$, the identity matrix, then $A$-statistical convergence coincides with ordinary convergence. The statistical convergence of various types of operators has been studied by several researchers (see mishra2017statistical ; duman2008statistical ; gadjiev2002some ; duman2006statistical ; gupta2009statistical ; ispir2008statistical ; gupta2014convergence ; gupta2021modifications ; gupta2021higher ; mursaleen2009invariant ; gurhan2015durrmeyer ; garrancho2019general ; chandra2022approximation ; agrawal2022stancu ; maurya2018approximation ). Now, we prove weighted Korovkin theorem via $A$-statistical convergence.
Here, we recall the weighted Korovkin type approximation theorem for the A-statistical convergence was given by Duman and Orhan in duman2004statistical .
Theorem 2
duman2004statistical Let $A$ be a non-negative regular summability matrix and let $\bar{\rho}_{1}$; $\bar{\rho}_{2}$ weight functions such that
@@
Proof: From (duman2004statistical, , p. 191, Th. 3), it is sufficient to show that $st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(t^{i},\cdot\right)-x^{i}\|_{2}=0$, where $i=0,1,2$.
@@
$\|W_{n}^{(\beta)}\left(t^{2},\cdot\right)-x^{2}\|_{2}\leq\sup_{x\geq 0}\left(\dfrac{x}{n\beta(1+x^{2})}\right)\leq\dfrac{1}{n\beta}$
@@
$U_{1}:=\left\{n:\frac{1}{n\beta}\geq\dfrac{\epsilon}{2}\right\}$
@@
Using theorem 1, we get desired result. $\blacksquare$
Lemma 5
@@
$st_{A}-\lim_{n\to\infty}nW_{n}^{(\beta)}\left((t-x)^{4},x\right)=\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}$
@@
Proof: Note that
$W_{n}^{(\beta)}\left((t-x)^{4},x\right)$$\leq$ $\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{n\beta(\beta+1)(\beta+2)}+\frac{4x^{2}\beta+11}{n^{2}\beta(\beta+1)}+\frac{x}{n^{3}\beta}.$
@@
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}\right|\leq\dfrac{h_{1}(\beta)}{n}x^{2}+\dfrac{h_{2}(\beta)}{n^{2}}x$
@@
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}\right|\leq B\left\{\dfrac{1}{n}+\dfrac{1}{n^{2}}\right\}$ (4.2)
where $B=\max\{h_{1}(\beta)b^{2},h_{2}(\beta)b\}$. Now, for a given $\epsilon>0$, define the following sets:
$D$ ::=:= $\left\{n:\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\frac{6x^{3}\left(\beta^{2}+5\beta+7\right)}{\beta(\beta+1)(\beta+2)}\right|\geq\epsilon\right\}$
$D_{1}$ ::=:= $\left\{n:\dfrac{1}{n}\geq\dfrac{\epsilon}{2B}\right\}$
$D_{2}$ ::=:= $\left\{n:\dfrac{1}{n}\geq\sqrt{\dfrac{\epsilon}{2B}}\right\}$
Hence, by inequality (4.2), we see that $D\subset D_{1}\cup D_{2}$. Then for any $j\in\mathbb{N}$, we have
$\sum_{n\in D}a_{jn}\leq\sum_{n\in D_{1}}a_{jn}+\sum_{n\in D_{2}}a_{jn}.$ (4.3)
@@
Hence the result. $\blacksquare$
Theorem 4 (Statistical Voronovskaya-type theorem for the operator $W_{n}^{(\beta)}$)
@@
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)=\dfrac{x}{2\beta}f^{\prime\prime}(x)$
@@
Proof. Let $f,f^{\prime},f^{\prime\prime}\in E$ and $x\in[0,b]$. Define the function $\Phi_{x}$ by
@@
$\left|n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)-\dfrac{x}{\beta}f^{\prime\prime}(x)\right|$$\leq$ $n\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|$ (4.4)
@@
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)=\dfrac{x}{\beta}f^{\prime\prime}(x)$
uniformly with respect to $x\in[0,b]$. $\blacksquare$
5 $\lambda\gamma$-statistical convergence
Recently, $\lambda\gamma$-statistical convergence of linear positive operators drawn the attention of several researchers (see karaisa2017alphabeta ; bozma2022approximation ; karaisaa2017alphabeta ; gogoi2021weighted ; karakaya2015korovkin ; braha2020korovkin ; abay2018alphabeta ; aktuuglu2016weighted ). Let $K$ be a set of natural numbers, then the density of $K$ is denoted by $\delta(K)$ and defined as
$\delta(K)=\lim_{n\to\infty}\dfrac{|K_{n}|}{n}$ (5.1)
where $K_{n}:=\{k\leq n:k\in K\}$ and $|S|$ represents the cardinality of $S$. Using the density function given in (5.1), Now, we introduced the concept of statistical convergence defined as fast1951convergence . A sequence $x_{n}$ is called statistically convergent to $l$ and denoted by $st-\lim_{n\to\infty}x_{n}=l$ if for every $\epsilon>0$, $\delta(K_{\epsilon})=0$ where $K_{\epsilon}:=\{k\leq n:|x_{k}-l|\geq\epsilon\}$.
Recently, the concept of $\lambda\gamma$-statistical convergence introduced by aktuuglu2014korovkin as follows:
Assume that $\lambda(n)$ and $\gamma(n)$ are two sequences of positive numbers satisfying the following conditions:
  1. $(C_{1})$
    $\lambda$ and $\gamma$ are both non-decreasing,
  2. $(C_{2})$
    $\gamma(n)\geq\lambda(n)$ for all $n\in\mathbb{N}$,
  3. $(C_{3})$
    $\gamma(n)-\lambda(n)\to\infty$ as $n\to\infty$,
and $\Lambda$ denotes the set of pairs $(\lambda,\gamma)$ satisfying $(C_{1})$, $(C_{2})$ and $(C_{3})$.
For each pair $(\lambda,\gamma)\in\Lambda$, and $K\subset\mathbb{N}$, we define $\delta^{\lambda,\gamma}(K)$ in the following way,
$\delta^{\lambda,\gamma}(K)=\lim_{n\to\infty}\dfrac{|K\cap I_{n}^{\lambda,\gamma}|}{\left(\gamma(n)-\lambda(n)+1\right)},$ (5.2)
where $I_{n}^{\lambda,\gamma}=[\lambda(n),\gamma(n)]$. Because of (5.2), we can state the following lemma (aktuuglu2014korovkin ).
Lemma 6
Let $K$ and $M$ be two subsets of $\mathbb{N}$ and $(\lambda,\gamma)\in\Lambda$. Then the following properties hold true.
  1. (i)
    $\delta^{\lambda,\gamma}(\phi)=0$,
  2. (ii)
    $\delta^{\lambda,\gamma}(\mathbb{N})=0$,
  3. (iii)
    If $K$ is a finite set, then $\delta^{\lambda,\gamma}(K)=0$,
  4. (iv)
    If $K\subset M$, then $\delta^{\lambda,\gamma}(K)\leq\delta^{\lambda,\gamma}(M)$,
  5. (v)
    $\delta^{\lambda,\gamma}(K\cup M)\leq\delta^{\lambda,\gamma}(K)+\delta^{\lambda,\gamma}(M).$
The following definition includes ordinary convergence but along with some well-known matrix methods such as statistical, and $\lambda$-statistical convergence (see aktuuglu2014korovkin ).
Definition 1
A sequence $x$ is said to be $\lambda\gamma$-statistically convergence to $l$ and denoted by $st_{\lambda\gamma}-\lim_{n\to\infty}x_{n}=l$, if for every $\epsilon>0,$
$\delta^{\lambda,\gamma}\left(\{k\in I_{n}^{\lambda,\gamma}:|x_{n}-l|\geq\epsilon\}\right)=\lim_{n\to\infty}\dfrac{\left|\{k\in I_{n}^{\lambda,\gamma}:|x_{n}-l|\geq\epsilon\}\right|}{\left(\gamma(n)-\lambda(n)+1\right)}=0.$
Remark 1
Take $\lambda(n)=1$ and $\gamma(n)=n$. Then $I_{n}^{\lambda,\gamma}=[1,n]$ and
$\delta^{\lambda,\gamma}\left(\{k\in[1,n]:|x_{n}-l|\geq\epsilon\}\right)=\lim_{n\to\infty}\dfrac{\left|\{k\leq n:|x_{n}-l|\geq\epsilon\}\right|}{n}=0.$
Thus for $\lambda(n)=1$ and $\gamma(n)=n$, $\lambda\gamma$-statistical convergence reduces to statistical convergence.
The following two examples show that $\lambda\gamma$-statistical convergence is a non-trivial extension of both ordinary and statistical convergence depending on the choice of $\lambda$ and $\gamma$.
Example 1
Consider the sequence
$c_{n}=\left\{\begin{array}[]{cc}n^{2},&\textrm{ if }n=m^{2}\textrm{ for some }m\in\mathbb{N},\\ \dfrac{1}{n},&\textrm{ otherwise,}\end{array}\right.$ (5.3)
taking $\lambda(n)=1$ and $\gamma(n)=n$, then $\lambda\gamma$-statistical convergence coincide with statistical convergence therefore $st_{\lambda\gamma}-\lim_{n\to\infty}c_{n}=0$.
Example 2
Take $\lambda(n)=2^{2n-1}$, $\gamma(n)=2^{2n}$ and let
$d_{n}=\left\{\begin{array}[]{cc}0,&\textrm{ if }n\in[2^{2m-1},2^{2m}-1]\textrm{ for }m\in\mathbb{N},\\ 1,&\textrm{ otherwise,}\end{array}\right.$ (5.4)
Then the sequence $(d_{n})$ does not converge in statistical sense. Since
|$\dfrac{\left|\{k\in I_{n}^{\lambda,\gamma}:|d_{k}|\geq\epsilon\}\right|}{\left(\gamma(n)-\lambda(n)+1\right)}\leq\dfrac{1}{\left(\gamma(n)-\lambda(n)+1\right)},$
we have $st_{\lambda\gamma}-\lim_{n\to\infty}d_{n}=0$.
Theorem 5
Let $(\lambda,\gamma)\in\Lambda$ and $\mu>0$ and $\beta>0$ be fixed, then
$st_{\lambda\gamma}-\lim_{n}\|W_{n}^{(\beta)}\left(f,\cdot\right)-f\|_{C[0,B]}=0$
holds for every $f\in E$.
Proof: By Lemma 2, we have $\|W_{n}^{(\beta)}\left(1,\cdot\right)-1\|_{C[0,B]}=0$ and $st_{\lambda\gamma}-\|W_{n}^{(\beta)}\left(t,\cdot\right)-t\|_{C[0,B]}=0$.
Again by Lemma 2
$\left\|W_{n}^{(\beta)}(t^{2},\cdot)-t^{2}\right\|_{C[0,B]}\leq\dfrac{x}{n\beta}.$
Now, let $\beta>0$ be a fixed number. For a given $s>0$ choose $\epsilon>0$ such that $\epsilon<s$ and define the following sets:
$V$ ::=:= $\{k\in I_{n}^{\lambda,\gamma}:\|W_{n}^{(\beta)}\left(f,\cdot\right)-f(\cdot)\|_{C[0,B]}\geq s\},$
$V_{1}$ ::=:= $\left\{k\in I_{n}^{\lambda,\gamma}:\dfrac{1}{k}\geq\dfrac{s\beta}{2B}\right\}.$
It is obvious that $V\subset V_{1}$. Therefore, Lemma 6 implies that,
$\delta^{\lambda,\gamma}(V)\leq\delta^{\lambda,\gamma}(V_{1}).$
Using $st_{\lambda\gamma}-\dfrac{1}{n}=0$, we have
$\delta^{\lambda\gamma}(V)=0.$
Hence using (aktuuglu2014korovkin, , Theorem 2), the proof is completed. $\blacksquare$
Now, we consider the following function spaces:
$B_{\rho}:=\{f:\mathbb{R}\to A\subset\mathbb{R}:|f(x)|\leq M_{f}\rho(x)\}$
$C_{\rho}:=\{f:\mathbb{R}\to A\subset\mathbb{R}:f\textrm{ is continuous over }\mathbb{R}\},$
where $\rho(x)$ is a weight function (i.e. $\rho(x)$ is continuous on $\mathbb{R}$ with $\lim_{|x|\to\infty}\rho(x)=\infty$ and $\rho(x)\geq 1$ for all $x\in\mathbb{R}$) and $M_{f}$ is a positive constant depending on $f$. Note that these spaces are Banach spaces endowed with the norm
$\|f\|_{\rho}=\sup_{x\in\mathbb{R}}\dfrac{|f(x)|}{\rho(x)}.$
We will use the following theorem.
Theorem 6 (Duman and Orhan duman2004statistical )
Let $(\lambda,\gamma)\in\Lambda$ and let $\rho_{1}$ and $\rho_{2}$ be weight functions satisfying $\lim_{|x|\to\infty}\dfrac{\rho_{1}(x)}{\rho_{2}(x)}=0$. Assume that $T_{n}$ is a sequence of linear positive operators acting from $C_{\rho_{1}}$ to $B_{\rho_{2}}$. Then for all $f\in C_{\rho_{1}}$,
$st_{\lambda\gamma}-\lim_{n}\|T_{n}(f,\cdot)-f\|_{\rho_{2}}=0$
if and only if
$st_{\lambda\gamma}-\lim_{n}\|T_{n}(F_{v},\cdot)-F_{v}\|_{\rho_{1}}=0,$
where $F_{v}(x)=\dfrac{x^{v}\rho_{1}(x)}{1+x^{2}}(v=0,1,2)$
Choosing the weight functions $\rho_{1}(x)=1+x^{2},\rho_{2}(x)=1+x^{2+\eta}$ for $\eta>0$ and $x\in[0,\infty)$.
Lemma 7
Let $\beta>0$ be fixed, then there exists a constant $N(\beta)>0$ such that
$\dfrac{1}{\rho_{2}(x)}W_{n}^{(\beta)}\left(\rho_{1}(x),x\right)\leq N(\beta)$
holds for all $x\in[0,\infty)$ and $n\in\mathbb{N}$. Furthermore, for all $f\in C_{\rho_{1}}$, we have
$\|W_{n}^{(\beta)}\left(f,\cdot\right)\|\leq N(\beta)\|f\|_{\rho_{1}}.$
Proof: By Lemma 2, we have
$\dfrac{1}{\rho_{2}(x)}W_{n}^{(\beta)}\left(\rho_{1}(t),x\right)$ $=$ $\dfrac{1}{1+x^{2+\mu}}\left[W_{n}^{(\beta)}\left(1,x\right)+W_{n}^{(\beta)}\left(t^{2},\cdot\right)\right]$
$\leq$ $\dfrac{1}{1+x^{2+\mu}}\left[1+x^{2}+\frac{1}{n\beta}x\right]$
$\leq$ $N(\beta).$
To prove the second inequality, we write
$\dfrac{1}{\rho_{2}(x)}\left|W_{n}^{(\beta)}\left(f,x\right)\right|=\dfrac{1}{\rho_{2}(x)}\left|W_{n}^{(\beta)}\left(\rho_{1}\dfrac{f}{\rho_{1}},x\right)\right|\leq\dfrac{\|f\|_{\rho_{1}}}{\rho_{2}(x)}W_{n}^{(\beta)}\left(\rho_{1}(x),x\right)\leq N(\beta)\|f\|_{\rho_{1}}.$ (5.5)
Taking supremum over $x\in[0,\infty)$ in the above inequality, immediately leads to the result. $\blacksquare$
The second inequality in the above lemma shows that $S^{(\beta)}_{n}:C_{\rho_{1}}\to B_{\rho_{2}}$.
Now, we will discussed following weighted $\lambda\beta$-statistical approximation theorem for the Wright operators (1.2):
Theorem 7
Let $(\lambda,\gamma)\in\Lambda$ and $\beta>0$ be fixed, then for each $f\in C_{1+x^{2}}[0,\infty)$ and $\eta>0$, we have
$st_{\lambda\gamma}-\lim_{n}\|W^{(\beta)}_{n}\left(f,\cdot\right)-f\|_{1+x^{2+\gamma}}=0.$
Proof: From Lemma 2, we have
$\|W^{(\beta)}_{n}\left(1,\cdot\right)-1\|_{1+x^{2}}=0.$
Using Lemma 2, we get
$st_{\lambda\gamma}-\|W^{(\beta)}_{n}\left(t,\cdot\right)-x\|_{1+x^{2}}=0.$
Again using Lemma 2, we have
$st_{\lambda\gamma}-\|W^{(\beta)}_{n}\left(t^{2},\cdot\right)-x^{2}\|_{1+x^{2}}$ $=$ $\sup_{x\in[0,\infty)}\dfrac{1}{1+x^{2}}\left[\dfrac{x}{n\beta}\right].$
For a given $r>0$, choose $\epsilon>0$ such that $\epsilon<r$. For fixed $\beta>0$, define sets:
$W$ ::=:= $\{k\in I_{n}^{\lambda,\gamma}:\|W_{n}^{(\beta)}\left(t^{2},\cdot\right)-x^{2}\|_{1+x^{2}}\geq r\}$
$W_{1}$ ::=:= $\left\{k\in I_{n}^{\lambda,\gamma}:\dfrac{1}{k}\geq\dfrac{r\beta}{2}\right\}.$
Since $W\subset W_{1}$, then
$\delta^{\lambda,\gamma}(W)\leq\delta^{\lambda,\gamma}(W_{1}).$
Using $st_{\lambda\gamma}-\dfrac{1}{n}=0$, we have
$\delta^{\lambda,\gamma}(W)=0.$
Finally, applying Theorem 6, hence the proof. $\blacksquare$
Remark 2
As per mention in Remark 1, if $\lambda(n)=1$ and $\gamma(n)=n$, $\lambda\gamma$-statistical convergence reduces to statistical convergence. Therefore, all results obtained here are also valid for statistical convergence.
Remark 3
Let $x_{n}$ be a non-decreasing sequence of positive numbers tending to $\infty$ such that
$x_{n+1}\leq x_{n}+1,\textrm{ and }x_{1}=1.$
Choosing $\lambda(n)=n-x_{n}+1$ and $\gamma(n)=n$, then clearly $I_{n}^{\lambda,\gamma}=[n-x_{n}+1,n]$. Moreover,
$\delta^{\lambda\gamma}\left(\{k:|y_{n}-L|\geq\epsilon\}\right)=\lim_{n\to\infty}\dfrac{\left|\{k\in[n-x_{n}+1,n]:|y_{k}-L|\geq\epsilon\}\right|}{x_{n}}$
this shows that for $\lambda(n)=n-x_{n}+1$ and $\gamma(n)=n$, $\lambda\gamma$-statistical convergence. Therefore, if we take $\lambda(n)=n-x_{n}+1$ and $\gamma_{n}(n)=n$ then all results obtained here are also valid in $x$-statistical sense.
6 Extensions
To obtain an approximation process in the space of integrable functions, we introduce the following new integral modification of Wright operators. Operators in the integral form are more suitable for approximating $L_{p}$ functions with $1\leq p<\infty$. The results for approximating $L_{p}$ functions may have potential applications in machine learning generalization analysis. For the real-valued bounded function $f$ defined on $[0,\infty)$ provided integral exists, we modified the operators (1.2) as follows:
Wright Operators involving Gamma Function:
The generalization of Wright operators involving gamma function is defined as
[…]
Wright Operators involving generalized Gamma Function:
The generalization of Wright operators involving generalized gamma function is defined as
[…]
The approximation properties of operators (6.1) and (6.2) are discussed in patel2024Wright ; patel2024Wright1 .
Kantorovich-type operators:
The Kantorovich-type operators is defined as follows
[…]
Durrmeyer type Modification:
The Durrmeyer type modification of Bernstein operators, Baskakov operators and Szasz–Mirakyan operators are studied in gupta2014convergence ; gupta2021higher ; gupta2021modifications ; ispir2008statistical . Motivated by such type of operators, we generalize the operators $W_{n}^{(\beta)}$ as follows:
[…]
where $b_{n,k}^{[d]}(t)=(-1)^{k}\dfrac{t^{k}}{k!}\xi_{n,d}^{(k)}(t)$. Here $\left\{\xi_{n,d}(x)\right\}$ is a sequence of functions, defined on $[0,b]$, $(b>0)$, and satisfying the following properties for every $n\in\mathbb{N}$ and $k\in\mathbb{N}\cup\{0\}$:
  • 1.
    $\xi_{n,d}\in C^{\infty}\left([a,b]\right),(0\leq a<b);$
  • 2.
    $\xi_{n,d}(0)=1$;
  • 3.
    $\xi_{n,d}$ is completely monotone, that is $(-1)^{k}\xi_{n,d}^{(k)}(x)\geq 0$, $(0\leq x\leq b)$
  • 4.
    there exists an integer $d$ such that $\xi_{n,d}^{(k+1)}(x)=-n\xi_{n+d,d}^{(k)}(x)$, $(n>\max\{0,-d\},x\in[0,b]).$
About these integral operators, we raise the problem to investigate their convergence in $L_{p}$-spaces.

### v2

@@ 3 Rate of Convergence
Proof.
From Lemma 2, we have
@@
$\leq$ $\dfrac{1}{1+x^{2}}\left[1+x^{2}+\dfrac{x(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\right]$
@@
Taking supremum over $x\in[0,\infty)$ in the above inequality, gives the result. ∎
@@
Theorem 1.
Let $\beta>1$, $f\in E$ and $\omega_{B+1}(f,\delta)$, $(B>0)$ be its modulus of continuity on the finite interval $0,B+1]\subset[0,\infty)$, then
@@
where $\delta_{n}(\beta,B)=\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}$ and $M_{f}(\beta,B)$ is an absolute constant depending on $f$, $\beta$ and $B$.
Proof.
Let $\beta>0$ be fixed. For $x\in[0,B]$ and $t\leq B+1$, we have following well-known inequality
@@
where $\delta>0$. Now, for $x\in[0,B]$ and $t>B+1$, using the fact that tx>1t-x>1, we have
@@
$\leq$ $6A_{f}(1+B^{2})\left(\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\right)+\left(1+\dfrac{\left[\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\right]^{1/2}}{\delta}\right)\omega_{B+1}(f,\delta)$
@@
where $M_{f}(\beta,B)=6A_{f}(1+B^{2})$ and $\delta_{n}(\beta,B)=\dfrac{B(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}$. Hence the proof. ∎
@@ 4 $A$-statistical Convergence
provided the series converges for each $n$. AA is said to be regular if $\lim_{n}(Ax)_{n}=L$ whenever $\lim_{n}x_{n}=L$. We say that, the sequence $x=(x_{n})$ is $A$-statistically convergent to $L$ and write $st_{A}-\lim_{n}x_{n}=L$ if for every $\epsilon>0$, $\lim_{n}\sum_{k:|x_{k}-L|\geq\epsilon}a_{nk}=0$.
Replacing $A$ by $C_{1}$, the Cesáro matrix of order one, the $A$-statistical convergence reduces to statistical convergence. Similarly, if we take $A=I$, the identity matrix, then $A$-statistical convergence coincides with ordinary convergence. The statistical convergence of various types of operators has been studied by several researchers (see [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]). Now, we prove weighted Korovkin theorem via $A$-statistical convergence.
Here, we recall the weighted Korovkin type approximation theorem for the A-statistical convergence was given by Duman and Orhan in [30].
Theorem 2.
[30] Let $A$ be a non-negative regular summability matrix and let $\bar{\rho}_{1}$; $\bar{\rho}_{2}$ weight functions such that
@@
Proof.
From [30, p. 191, Th. 3], it is sufficient to show that $st_{A}-\lim_{n}\|W_{n}^{(\beta)}\left(t^{i},\cdot\right)-x^{i}\|_{2}=0$, where $i=0,1,2$.
@@
$\|W_{n}^{(\beta)}\left(t^{2},\cdot\right)-x^{2}\|_{2}\leq\sup_{x\geq 0}\left(\dfrac{x^{2}(1+2\beta)}{\beta n(1+x^{2})}+\dfrac{x\beta^{2}}{n^{2}(1+x^{2})}\right)\leq\dfrac{1+2\beta}{n\beta}+\dfrac{\beta^{2}}{n^{2}}$
@@
$U_{1}:=\left\{n:\dfrac{(1+2\beta)}{\beta n}+\dfrac{\beta^{2}}{n^{2}}\geq\dfrac{\epsilon}{2}\right\}$
@@
Using theorem 1, we get desired result.∎
Lemma 5.
@@
$st_{A}-\lim_{n\to\infty}nW_{n}^{(\beta)}\left((t-x)^{4},x\right)=\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta.$
@@
Proof.
Note that
$W_{n}^{(\beta)}\left((t-x)^{4},x\right)$ $\leq$ $\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\frac{x^{3}}{n}+\left(\frac{6\beta^{2}+12\beta+7}{\beta(\beta+1)}+\frac{4}{\beta}+12+4\beta+6\beta^{2}\right)\frac{x^{2}}{n^{2}}$
$+\left(4\beta^{3}+4\beta^{2}+6\beta+4+\frac{1}{\beta}\right)\frac{x}{n^{3}}+\frac{\beta^{4}}{n^{4}}.$
@@
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\right|\leq\dfrac{h_{1}(\beta)}{n}x^{2}+\dfrac{h_{2}(\beta)}{n^{2}}x+\dfrac{\beta^{4}}{n^{3}}$
@@
$\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\right|\leq B\left\{\dfrac{1}{n}+\dfrac{1}{n^{2}}+\dfrac{1}{n^{3}}\right\}$ (4.2)
where $B=\max\{h_{1}(\beta)b^{2},h_{2}(\beta)b,\beta^{4}\}$. Now, for a given $\epsilon>0$, define the following sets:
$D$ $:=$ $\left\{n:\left|nW_{n}^{(\beta)}\left((t-x)^{4},x\right)-\left(\frac{4\beta+6}{\beta(\beta+1)(\beta+2)}+\frac{18}{\beta}+12+4\beta\right)\right|\geq\epsilon\right\}$
$D_{1}$ $:=$ $\left\{n:\dfrac{1}{n}\geq\dfrac{\epsilon}{2B}\right\}$
$D_{2}$ $:=$ $\left\{n:\dfrac{1}{n}\geq\sqrt{\dfrac{\epsilon}{2B}}\right\}$
$D_{3}$ $:=$ $\left\{n:\dfrac{1}{n}\geq\sqrt[3]{\dfrac{\epsilon}{B}}\right\}$
Hence, by inequality (4.2), we see that $D\subset D_{1}\cup D_{2}\cup D_{3}$. Then for any $j\in\mathbb{N}$, we have
$\sum_{n\in D}a_{jn}\leq\sum_{n\in D_{1}}a_{jn}+\sum_{n\in D_{2}}a_{jn}\sum_{n\in D_{3}}a_{jn}.$ (4.3)
@@
Hence the result. ∎
Theorem 4 (Statistical Voronovskaya-type theorem for the operator $W_{n}^{(\beta)}$).
@@
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)=\frac{1+2\beta+2\beta^{2}}{\beta}\dfrac{x}{2}f^{\prime\prime}(x)$
@@
Proof.
Let $f,f^{\prime},f^{\prime\prime}\in E$ and $x\in[0,b]$. Define the function $\Phi_{x}$ by
@@
$\left|n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)-\frac{1+2\beta+2\beta^{2}}{2\beta}\,xf^{\prime\prime}(x)\right|$ $\leq$ $n\left|W_{n}^{(\beta)}((t-x)^{2}\Phi_{x}(t),x)\right|$ (4.4)
@@
$st_{A}-\lim_{n\to\infty}n\left(W_{n}^{(\beta)}(f,x)-f(x)\right)\frac{1+2\beta+2\beta^{2}}{2\beta}\,xf^{\prime\prime}(x)$
uniformly with respect to $x\in[0,b]$. ∎